# Adapting Physics-Informed Neural Networks To Optimize ODEs in Mosquito Population Dynamics

### v1

@@ 6 Discussion and Future Work
Our research using PINNs revealed certain limitations. In the inverse problem setup, the PINN tended to favor solutions with smaller gradients that still met the data and ODE constraints. Furthermore, the capability of the framework to extrapolate in inverse problems remains uncertain; with lack of appropriate volume of data, the networks may converge to arbitrary solutions. This aspect underscores the need for incorporating additional regularization techniques or constraints in order to guide the network towards more plausible solutions. In the current implementation, the models consider only time as an input, without accounting for external factors. This restriction significantly limits predictive performance in real-world applications, as the coordination inputs and the encoded physical laws may not fully capture the underlying mechanisms of the processes. Allowing for external variables would not only enhance performance but also add flexibility to the framework. Although the framework has shown promising results on test systems, its scalability and ability to generalize to other types of dynamical systems have not been thoroughly evaluated. Our current work focuses on more extensive testing on a wider range of ODE systems as this is seen as essential to further develop the capabilities of PINN systems.
Acknowledgement
This work was supported by Science Foundation Ireland through the Insight Centre for Data Analytics (SFI/12/RC/2289_P2) and by COST Action CA20108, supported by COST (European Cooperation in Science and Technology).
Appendix A Error Metrics
@@
Appendix B Inverse Problem of Lorenz system
Shown in figure 7
Figure 7: OdePINN framework solves inverse problem with Lorenz system, $U$ approximation of the solution.
Appendix C Mosquito ODE system parameters
Shown in table 4.
Parameter Description Value Unit
$\tau$ Temperature ${}^{\circ}C$
$\gamma_{Aem}$ Development rate of emerging adults 1.143 $days^{-1}$
$\gamma_{Ab}$ Development rate of bloodseeking adults 0.885 $days^{-1}$
$\gamma_{Ao}$ Ovipositing adult development rate 2 $days^{-1}$
$f_{E}(>0)$ Egg development rate $0.16\cdot\left(e^{\left[0.105(\tau-10)\right]}-e^{\left[0.105(38-10)-\frac{1}{5.007}(38-\tau)\right]}\right)$ $days^{-1}$
$f_{P}$ Pupa development rate $0.021\cdot\left(e^{\left[{0.162(\tau-10)}\right]}-e^{\left[{0.162(38-10)-\frac{1}{5.007}(38-\tau)}\right])}\right)$ $days^{-1}$
$f_{L}$ Larva development rate $f_{P}/1.65$ $days^{-1}$
$f_{Ag}(>0)$ Development rate of gestating adults $\frac{\tau-9.8}{64.4}$ $days^{-1}$
$m_{E}$ Egg mortality rate $m_{E}=\mu_{E}$ $days^{-1}$
$m_{L}$ Larval mortality rate $\exp\left[{-\tau/2}\right]+\mu_{L}$ $days^{-1}$
$m_{P}$ Pupa mortality rate $\exp\left[{-\tau/2}\right]+\mu_{P}$ $days^{-1}$
$m_{A}(>\mu_{A})$ Mortality rate of $Ab$, $-0.005941+0.002965\cdot\tau$ $days^{-1}$
$\mu_{E}$ Minimum egg mortality rate 00 $days^{-1}$
$\mu_{L}$ Minimum larval mortality rate $0.0304$ $days^{-1}$
$\mu_{P}$ Minimum pupa mortality rate $0.0146$ $days^{-1}$
$\mu_{em}$ Mortality rate during emergence $0.1$ $days^{-1}$
$\mu_{r}$ Mortality rate during bloodseeking $0.08$ $days^{-1}$
$\mu_{A}$ Minimum adult mortality rate $\frac{1}{43}$ $days^{-1}$
$\kappa_{L}$ Carrying capacity for larvae $8\cdot 10^{8}$ $days^{-1}$
$\kappa_{P}$ Carrying capacity for pupae $10^{7}$ $days^{-1}$
$\sigma$ Sex ratio at emergence 0.5 -
$\beta$ Number of eggs per $Ao$ $\beta_{1}=141(np),\beta_{2}=80(p)$
@@ Appendix C Mosquito ODE system parameters
Appendix D Structural identifiability of parameters
Shown in figure 8
Figure 8: Structural identifiability of parameters over time in the Mosquito inverse problem, the sine-shape temperature. This is achieved by expressing the system as a system of linear equations, the parameters as unknown variables, and analyzing the Reduced row-echelon form of the coefficient matrix, performed separately for each time $t$. In the figure, identifiable parameters are defined as free parameters which can get arbitrary values. The values are rounded to 6-digit precision, aligning with the PINN’s level of precision after training. The figure explains the PINN’s inaccuracies shown in figure 6.

### v2

@@ Methods
Our research using PINNs revealed certain limitations. In the inverse problem setup, the PINN tended to favor solutions with smaller gradients that still met the data and ODE constraints. Furthermore, the capability of the framework to extrapolate in inverse problems remains uncertain; with lack of appropriate volume of data, the networks may converge to arbitrary solutions. This aspect underscores the need for incorporating additional regularization techniques or constraints in order to guide the network towards more plausible solutions. In the current implementation, the models consider only time as an input, without accounting for external factors. This restriction significantly limits predictive performance in real-world applications, as the coordination inputs and the encoded physical laws may not fully capture the underlying mechanisms of the processes. Allowing for external variables would not only enhance performance but also add flexibility to the framework. Although the framework has shown promising results on test systems, its scalability and ability to generalize to other types of dynamical systems have not been thoroughly evaluated. Our current focus is on the development of a new PINN framework that encodes complex, learned interactions of air temperature, precipitation and relative humidity, to improve and demonstrate the effectiveness of this approach in addressing inverse problems, such as inferring mosquito development and mortality rates. Accurate mosquito population dynamics modelling is essential for predicting the outbreaks of arbovirus diseases. By improving predictive accuracy, PINNs can provide more reliable insights into population trends and help design targeted control measures (e.g., optimal times for pesticide application). This leads to more effective, data-informed vector control strategies that can be adapted across different geographic and climatic contexts.
Supporting information
S1 Appendix.
Error metrics.
@@
S1 Fig.
OdePINN framework solves inverse problem with Lorenz system, $U$ approximation of the solution (Fig 9).
S1 Table.
ODE Parameters (Table 4).
S2 Fig.
Model scheme showing the 10 develpmental stages within the mosquito life cycle (Fig 10): Egg ($E$), Larva ($L$), Pupa ($P$), Emerging Adults ($A_{em}$), Nulliparous Bloodseeking Adults ($A_{b1}$), Nulliparous Gestating Adults ($A_{g1}$), Nulliparous Ovipositing Adults ($A_{o1}$), Parous Bloodseeking Adults ($A_{b2}$), Parous Gestating Adults ($A_{g2}$) and Parous Ovipositing Adults ($A_{o2}$) (source: [11]).
S3 Fig.
Identifiability of parameters over time in the Mosquito inverse problem, the sine-shape temperature (Fig 11). This is achieved by expressing the system as a system of linear equations, the parameters as unknown variables, and analyzing the Reduced row-echelon form of the coefficient matrix, performed separately for each time $t$. In the figure, identifiable parameters are defined as free parameters which can get arbitrary values. The values are rounded to 6-digit precision, aligning with the PINN’s level of precision after training. The figure explains the PINN’s inaccuracies shown in Fig 8.
Fig 9: OdePINN framework solves inverse problem with Lorenz system, $U$ approximation of the solution.
{adjustwidth}
-2.00in0in Parameter Description Value Unit $\tau$ Temperature ${}^{\circ}C$ $\gamma_{Aem}$ Development rate of emerging adults 1.143 $days^{-1}$ $\gamma_{Ab}$ Development rate of bloodseeking adults 0.885 $days^{-1}$ $\gamma_{Ao}$ Ovipositing adult development rate 2 $days^{-1}$ $f_{E}(>0)$ Egg development rate $0.16\cdot\left(e^{\left[0.105(\tau-10)\right]}-e^{\left[0.105(38-10)-\frac{1}{5.007}(38-\tau)\right]}\right)$ $days^{-1}$ $f_{P}$ Pupa development rate $0.021\cdot\left(e^{\left[{0.162(\tau-10)}\right]}-e^{\left[{0.162(38-10)-\frac{1}{5.007}(38-\tau)}\right])}\right)$ $days^{-1}$ $f_{L}$ Larva development rate $f_{P}/1.65$ $days^{-1}$ $f_{Ag}(>0)$ Development rate of gestating adults $\frac{\tau-9.8}{64.4}$ $days^{-1}$ $m_{E}$ Egg mortality rate $m_{E}=\mu_{E}$ $days^{-1}$ $m_{L}$ Larval mortality rate $\exp\left[{-\tau/2}\right]+\mu_{L}$ $days^{-1}$ $m_{P}$ Pupa mortality rate $\exp\left[{-\tau/2}\right]+\mu_{P}$ $days^{-1}$ $m_{A}(>\mu_{A})$ Mortality rate of $Ab$, $-0.005941+0.002965\cdot\tau$ $days^{-1}$ $\mu_{E}$ Minimum egg mortality rate 00 $days^{-1}$ $\mu_{L}$ Minimum larval mortality rate $0.0304$ $days^{-1}$ $\mu_{P}$ Minimum pupa mortality rate $0.0146$ $days^{-1}$ $\mu_{em}$ Mortality rate during emergence $0.1$ $days^{-1}$ $\mu_{r}$ Mortality rate during bloodseeking $0.08$ $days^{-1}$ $\mu_{A}$ Minimum adult mortality rate $\frac{1}{43}$ $days^{-1}$ $\kappa_{L}$ Carrying capacity for larvae $8\cdot 10^{8}$ $days^{-1}$ $\kappa_{P}$ Carrying capacity for pupae $10^{7}$ $days^{-1}$ $\sigma$ Sex ratio at emergence 0.5 - $\beta$ Number of eggs per $Ao$ $\beta_{1}=141(np),\beta_{2}=80(p)$
@@
Fig 10: Model scheme showing the 10 develpmental stages within the mosquito life cycle: Egg ($E$), Larva ($L$), Pupa ($P$), Emerging Adults ($A_{em}$), Nulliparous Bloodseeking Adults ($A_{b1}$), Nulliparous Gestating Adults ($A_{g1}$), Nulliparous Ovipositing Adults ($A_{o1}$), Parous Bloodseeking Adults ($A_{b2}$), Parous Gestating Adults ($A_{g2}$) and Parous Ovipositing Adults ($A_{o2}$) (source: [11])
Fig 11: Identifiability of parameters over time in the Mosquito inverse problem, the sine-shape temperature. This is achieved by expressing the system as a system of linear equations, the parameters as unknown variables, and analyzing the Reduced row-echelon form of the coefficient matrix, performed separately for each time $t$. In the figure, identifiable parameters are defined as free parameters which can get arbitrary values. The values are rounded to 6-digit precision, aligning with the PINN’s level of precision after training. The figure explains the PINN’s inaccuracies shown in Fig 8.